OpenStudy (anonymous):

HELP PLEASE!!!! The Ward family is redecorating several rooms of their house. They decided to use carpet tiles in the family room. The room has an area of 176 square feet and is 5 feet longer than it is wide. Find the dimensions of the family room.

OpenStudy (anonymous):

11by16

OpenStudy (anonymous):

THANK YOU!!

OpenStudy (anonymous):

all you had to do is divid 176 by number till you get 2 numbers and 1 is 5 greater than the other

OpenStudy (anonymous):

You can set-up a system of equations. If we assume the room is a rectangle since the length is longer than the width. Area of a rectangle: \[Area = length \times width\] Length (L) = W +5 Width = W Area = 176 176 = L x W Substituting what we know. 176 = (W+5) x W You can solve for the Width, W , which in turn will solve for the Length, L.

OpenStudy (anonymous):

like my answer become a fan of me

OpenStudy (anonymous):

you can also message me when you need help

OpenStudy (anonymous):

It is against the Code of Conduct to give answers and I do not want to steal the education of others. Code of Conduct: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

i told her what she could do to get the answer after and lots of people give other people answers